Technische Universität Berlin Sommersemester 2006 Fakultät II – Institut für Mathematik Dr. M. Ehrhardt

Projektblatt Gletscher zur Vorlesung “Theorie und Numerik hyperbolischer Erhaltungsgleichungen”

Literatur:

(N1) J.F. Nye, The response of glaciers and ice-sheets to seasonal and climatic changes, Proc. Roy. Soc. A 256 (1960), 559–584.

(N2) J.F. Nye, The response of glaciers to changes in the rate of nourishment and wastage, Proc. Roy. Soc. A 275 (1963), 87–112.

(W) G. Whitham, Linear and Nonlinear Waves, Wiley-Interscience, 1974, Kapitel 3.3.